OpenStudy (anonymous):

(a) Find the coordinates of the points of intersection of y= 4/x and 2x+y=9. Answer= (1/2,8) and (4,1) ~~I know how to get this answer~~ (b) Sketch both graphs for values of x such that x>0. ~~ Need help ~~ (c) Calculate the area between the graphs. ~~ Need help~~

OpenStudy (anonymous):

the answer to part c is :\[\int\limits_{1/2}^{4}4/x +2x-9\] = \[-4\ln x+x^{2}-9x\] from 1/2 to 4

OpenStudy (anonymous):

for understanding why I used this way you can plot the graph and see the area between them...=)(I hope it be useful )

OpenStudy (anonymous):

I should have written it this way : \[\int\limits_{1/2}^{4}4/x - (9 - 2x)\]

OpenStudy (anonymous):

no sorry I should have written it this way ,\[\int\limits_{1/2}^{4}-2x +9 - 4/x\]

OpenStudy (anonymous):

15 3/4 - 12In2 is the answer. How do you use substitution in order to get to that answer?

OpenStudy (anonymous):

ok the answer to the above integral I wrote is :\[-x ^{2 }+9x -4\ln x\]from 1/2 to 4 then We have : \[(-4)^{2}+9(4)-4ln 4+1/4-9/2-4\ln 2\]

OpenStudy (anonymous):

sorry \[-(4)^{2}\] is right for the first part

OpenStudy (anonymous):

\[4\ln 4\]= \[8\ln 2\]

OpenStudy (anonymous):

The whole answer is 15 3/4 - 12In2 ?

OpenStudy (anonymous):

I think is 53/4 - 12ln2

OpenStudy (anonymous):

isn't it right?

OpenStudy (anonymous):

but why? That is the answer... yes :) but I don't know how to get it.

OpenStudy (anonymous):

let me som min I try to explain (my problem as you've understand is my poor English=))

OpenStudy (anonymous):

Haha, not poor english! Ok. I'll let you explain.

OpenStudy (anonymous):

−x^2+9x−4lnx from 1/2 to 4 then We have : (-(4)^2+9(4)−4ln4)-(-1/4+9/2+4ln2)= -16+36 -17/4-8ln2-4ln2 = 20-17/4 -12ln2 = 53/4-12ln2 that's all

OpenStudy (anonymous):

Oh, Ok. Thank you!

OpenStudy (anonymous):

Don't mention it!=)